OpenStudy (dcolley):

Solve the rational inequality -x+7/x-9 >=0

OpenStudy (anonymous):

\[-x+7 >0\]\[-x>-7\]\[x<7\]

OpenStudy (dcolley):

in interval notatold be (-infinity,7] in union (infinity,9)

OpenStudy (anonymous):

\[ \frac{7-x}{x-9} \ge 0\]

OpenStudy (anonymous):

unknwn the denominator. Multiply by denomintor squared.

OpenStudy (anonymous):

\[(7-x)(x-9) \ge 0\]